OpenStudy (mikezack123):

Choose the correct x-intercept and y-intercept of the equation 3x + 2y = 6. x-intercept: (0, 2), y-intercept: (0, 3) x-intercept: (2, 0), y-intercept: (0, 3) x-intercept: (3, 0), y-intercept: (2, 0) x-intercept: (0, 3), y-intercept: (0, 2)

OpenStudy (marissalovescats):

You probably want to put it in y=mx+b form first

OpenStudy (anonymous):

There are a few ways you can do this. The quickest way for me would probably be to go through the answer choices, plugging in the points, and seeing if they make the equation true. If one or both of the points in an answer choice returns a false statement, that answer choice is out. You could also rearrange it into a different form such as y=mx+b.

OpenStudy (anonymous):

x-intercept: (2, 0), y-intercept: (0, 3) #300

OpenStudy (jdoe0001):

set x = 0, solve for "y", to get the x-intercept set y =0, solve for "x" to get the y-intercept

OpenStudy (mikezack123):

y=-3/2+6

OpenStudy (jdoe0001):

I think I got that backwards :/

OpenStudy (jdoe0001):

set x = 0, solve for "y", to get the y-intercept set y =0, solve for "x" to get the x-intercept

OpenStudy (mikezack123):

x-intercept: (2, 0), y-intercept: (0, 3)

OpenStudy (mikezack123):

?

OpenStudy (texaschic101):

set x = 0 to find the y intercept and set y = 0 to find the x intercept....it is that easy